# A Lightweight Library for Energy-Based Joint-Embedding Predictive Architectures

### v1

@@ 1 Introduction
JEPA builds on a rich history of self-supervised representation learning (Chen et al., 2020; He et al., 2020; Grill et al., 2020; Zbontar et al., 2021; Chen & He, 2021) while avoiding the need for negative samples. JEPAs have demonstrated strong performance for visual representation learning (Assran et al., 2023) and have been extended to video understanding (Bardes et al., 2024) and world modeling for planning (Assran et al., 2025; Sobal et al., 2025; Zhou et al., 2024a; Terver et al., 2026). Despite this growing body of work, accessible implementations that bridge theoretical principles and practical application remain scarce. Production-scale implementations are designed for large-scale training and are challenging to navigate. World model implementations like DINO-WM (Zhou et al., 2024a) and JEPA-WMs (Terver et al., 2026) enable planning on simple environments but rely on particular setups, e.g., frozen pre-trained encoders. As a result, while JEPAs have shown promises, they remain with a high barrier to entry–which we hope to address in this study. This paper introduces EB-JEPA, an open-source library that addresses this gap through modular, well-documented implementations of JEPA-based models trainable at small scale with simple, concise code designed for educational purposes and rapid experimentation. Our contributions are:
@@ 2 Related Work
EB-JEPA builds on the JEPA framework and non-contrastive self-supervised learning (Adrien Bardes, 2022; Zbontar et al., 2021; Grill et al., 2020; Chen & He, 2021; Oquab et al., 2024). Recent theoretical work has provided deeper understanding of these methods: Shwartz-Ziv et al. (2023) analyze VICReg from an information-theoretic perspective, while Balestriero & LeCun (2022) show connections between contrastive and non-contrastive methods and spectral embedding. While I-JEPA and V-JEPA focus on masked prediction within single images or videos, our action-conditioned example extends this to interactive settings where actions determine future states. Recent work has shown that JEPA-style pretraining leads to emergent understanding of intuitive physics (Garrido et al., 2025), motivating the use of such architectures for world modeling. Importantly, JEPAs differ fundamentally from reconstruction-based methods such as MAE (He et al., 2021) and VideoMAE (Tong et al., 2022; Wang et al., 2023), which predict in pixel space rather than representation space. Balestriero & Lecun (2024) provide theoretical analysis showing that reconstruction-based learning can produce uninformative features for perception, further motivating the joint-embedding paradigm that our library focuses on.
@@ 3 Preliminaries: A Unified JEPA Framework
The key challenge in training EBMs is preventing collapse: a degenerate solution where the energy is uniformly low for all inputs. Classical EBMs address this through contrastive methods that explicitly push up the energy of “negative” samples (Hinton, 2002; Chen et al., 2020; He et al., 2020). JEPAs instead rely on regularization-based approaches (Adrien Bardes, 2022; Zbontar et al., 2021; Balestriero & LeCun, 2025) that maintain representation diversity without requiring negative samples. This insight has proven powerful across self-supervised learning (Grathwohl et al., 2020; Grill et al., 2020). In the JEPA framework, we instantiate the energy principle by defining energy as prediction error in representation space. With the regularizer $\mathcal{R}$ and a given prediction loss $\mathcal{L}_{\text{pred}}$, the JEPA general training objective takes the form
@@
The key challenge in training JEPAs is preventing representation collapse, where the encoder learns trivial constant representations. EB-JEPA implements two regularization families. VICReg (Adrien Bardes, 2022) prevents collapse through two complementary terms. The variance term ensures each feature dimension has sufficient spread across the batch and reads
@@ 6 Future Directions
Our experiments highlight the critical role of regularization in preventing representation collapse, yet the theoretical understanding of why certain regularization combinations work remains incomplete. EB-JEPA provides a testbed for systematically studying regularization dynamics: investigating the interplay between variance, covariance, temporal similarity, and inverse dynamics terms (Adrien Bardes, 2022; Balestriero & LeCun, 2025; Sobal et al., 2022); understanding when each becomes necessary; and developing principled methods for automatic hyperparameter selection. The controlled, single-GPU setting enables rapid iteration on these fundamental questions without the confounding factors introduced by large-scale distributed training.

### v2

@@ 1 Introduction
JEPA builds on a rich history of self-supervised representation learning (Chen et al., 2020; He et al., 2020; Grill et al., 2020; Zbontar et al., 2021; Chen & He, 2021). JEPAs have demonstrated strong performance for visual representation learning (Assran et al., 2023) and have been extended to video understanding (Bardes et al., 2024) and world modeling for planning (Assran et al., 2025; Sobal et al., 2025; Zhou et al., 2024a; Terver et al., 2026). Despite this growing body of work, accessible implementations that bridge theoretical principles and practical application remain scarce. Production-scale implementations are designed for large-scale training and are challenging to navigate. World model implementations like DINO-WM (Zhou et al., 2024a) and JEPA-WMs (Terver et al., 2026) enable planning on simple environments but rely on particular setups, e.g., frozen pre-trained encoders. As a result, while JEPAs have shown promises, they remain with a high barrier to entry–which we hope to address in this study. This paper introduces EB-JEPA, an open-source library that addresses this gap through modular, well-documented implementations of JEPA-based models trainable at small scale with simple, concise code designed for educational purposes and rapid experimentation. Our contributions are:
@@ 2 Related Work
EB-JEPA builds on the JEPA framework (Assran et al., 2023; Bardes et al., 2024), focusing on their subclass using regularization-based collapse prevention (Bardes et al., 2022; Zbontar et al., 2021; Balestriero & LeCun, 2025) rather than stop-gradient techniques (Grill et al., 2020; Chen & He, 2021; Oquab et al., 2024). Recent theoretical work has provided deeper understanding of these methods: Shwartz-Ziv et al. (2023) analyze VICReg from an information-theoretic perspective, while Balestriero & LeCun (2022) show connections between contrastive and non-contrastive methods and spectral embedding. While I-JEPA and V-JEPA focus on masked prediction within single images or videos, our action-conditioned example extends this to interactive settings where actions determine future states. Recent work has shown that JEPA-style pretraining leads to emergent understanding of intuitive physics (Garrido et al., 2025), motivating the use of such architectures for world modeling. Importantly, JEPAs differ fundamentally from reconstruction-based methods such as MAE (He et al., 2021) and VideoMAE (Tong et al., 2022; Wang et al., 2023), which predict in pixel space rather than representation space. Balestriero & Lecun (2024) provide theoretical analysis showing that reconstruction-based learning can produce uninformative features for perception, further motivating the joint-embedding paradigm that our library focuses on.
@@ 3 Preliminaries: A Unified JEPA Framework
The key challenge in training EBMs is preventing collapse: a degenerate solution where the energy is uniformly low for all inputs. Different strategies address this challenge: contrastive methods push up energy on negative samples (Hinton, 2002; Chen et al., 2020; He et al., 2020); stop-gradient and exponential moving average (EMA) techniques break symmetry (Grill et al., 2020; Chen & He, 2021; Assran et al., 2023; Bardes et al., 2024); and regularization-based approaches maintain representation diversity without negative samples (Bardes et al., 2022; Zbontar et al., 2021; Balestriero & LeCun, 2025). Our library focuses on the regularization approach: we instantiate JEPAs with explicit regularization losses to prevent collapse, defining energy as prediction error in representation space. With the regularizer $\mathcal{R}$ and a given prediction loss $\mathcal{L}_{\text{pred}}$, the JEPA general training objective takes the form
@@
The key challenge in training JEPAs is preventing representation collapse, where the encoder learns trivial constant representations. EB-JEPA implements two regularization families. VICReg (Bardes et al., 2022) prevents collapse through two complementary terms. The variance term ensures each feature dimension has sufficient spread across the batch and reads
@@ 6 Future Directions
Our experiments highlight the critical role of regularization in preventing representation collapse, yet the theoretical understanding of why certain regularization combinations work remains incomplete. EB-JEPA provides a testbed for systematically studying regularization dynamics: investigating the interplay between variance, covariance, temporal similarity, and inverse dynamics terms (Bardes et al., 2022; Balestriero & LeCun, 2025; Sobal et al., 2022); understanding when each becomes necessary; and developing principled methods for automatic hyperparameter selection. The controlled, single-GPU setting enables rapid iteration on these fundamental questions without the confounding factors introduced by large-scale distributed training.